[m] = 3m if m is odd
[m] = 1/2m if m is even
[9]x[6] is equivalent to
a. 81
b. 54
c. 27
d. 18
e. 6 (can't remember)
answer is 27 but I don't have explanation from practice test

Answer by rapaljer(4667)

I'm sorry, but I get the answer of 81.

[m] = 3m if m is odd
[m] = 1/2m if m is even

Since 9 is odd, [9]=3*9=27
Since 6 is even, [6]= 1/2*6 = 3

[9]x[6] is equivalent to 27x3 = 81.

Maybe the answer is wrong. You may want to re-post this question and see if anyone else has a different idea about it.

R^2 Retired from SCC